OpenStudy (anonymous):

derivative of -x^2+x

OpenStudy (anonymous):

\[ax^n=anx^{n-1}\]where "a" is a constant, and n is any real number

OpenStudy (anonymous):

remember x = x^1 and x^0 = 1

OpenStudy (anonymous):

Now knowing this, what do you think the answer is?

OpenStudy (anonymous):

@ChmE meant to write (d/dx)(ax^n) = a n x^(n-1)